X/2 + x/3 >17/6


1
Expert's answer
2022-03-28T16:06:19-0400
"\frac{x}{2}+\frac{x}{3}>\frac{17}{6}"

Multiply both sides of the inequality by 6


"3x+2x>17"


"5x>17"

Now divide both sides of the inequality by 5

"x>\frac{17}{5}"